The following information is available for the 22,400 units of X Company's only product sold in 2019:

Selling price $64.80
Variable costs per unit $21.80
Total fixed costs $336,000

Management believes it can increase the selling price in 2020 with no change in unit sales. Variable costs per unit and fixed costs would remain the same. What would the selling price have to be in 2020 in order for X Company's 2020 profit to be $784,000?
Answer #1

Target contribution margin would be=Fixed cost+Target profit

=(336,000+784,000)=$1120000

Hence contribution margin per unit=1120000/22400

=$50 per unit

Contribution margin=Sales-Variable cost  

Hence sales price required=(50+21.8)

=$71.8 per unit